hs fegarding rates An American firm is under obligation to pay interests of Cans 1010000 and Cans 705000 on 31st July and 30th September respectively. The Firm is risk averse andi policy is to hedge the risks involved in all foreign currency transactions. The Finance Manager of the firm is thinking of hedging the risk considering two methods i.e. fixed forward or option contracts. It is now June 30. Following quotations regarding rates of exchange, USS per Cans, from the firms bank were obtained: Spot 1 Month Forward 3 Months Forward 0.9284-0.9288 0.9301 0.9356 Price for a CanS/USS option on a U.S. stock exchange (cents per CanS, payable on purchase of the option, contract size Can$ 50000) are as follows: Strike Price Calls Puts (US$/Cans) July Sept. July Sept. 0.93 1.56 2.56 0.88 1.75 0.94 1.02 NA NA NA 0.95 0.65 1.64 1.92 2.34 According to the suggestion of finance manager if options are to be used, one month option should be bought at a strike price of 94 cents and three month option at a strike price of 95 cents and for the remainder uncovered by the options the firm would bear the risk itself. For this, it would use forward rate as the best estimate of spot. Transaction costs are ignored. Recommend, which of the above two methods would be appropriate for the American fim to hedge its forelgn exchange risk on the two interest payments. t23 Sheet24 Sheeti7 Sheet18 Sheet19 Sheet20 Sheet21 Sheet22 Sheet?? s
Answer #1
Answer
Forward Market Cover
Hedge the risk by buying Can$ in 1 and 3 months time will be:
July - 1010000 X 0.9301 = US $ 939401
Sept. - 705000 X 0.9356 = US $ 659598
Option Contracts
July Payment = 1010000/ 50,000 = 20.20
September Payment = 705000/ 50,000 = 14.10
Company would like to take out 20 contracts for July and 14 contracts for September respectively. Therefore costs, if the options were exercised, will be:-
july sept
Can$ US $ Can $ US $
Covered by Contracts 1000000 940000 700000 665000
Balance bought at spot rate 10000 9301 5000 4678
Option Costs:
Can $ 50000 x 20 x 0.0102 10200 0
Can $ 50000 x 14 x 0.0164 0 11480
Total cost in US $ of using Option Contract
Decision: As the firm is stated as risk averse and the money due to be paid is certain, a fixed forward contract, being the cheapest alternative in the both the cases, would be recommended.